OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x+y=4 x-y=-6 A. (2, 2) B. (1, −7) C. (−1, 5) D. (−2, 4)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@karolynxx @hhmorris98

OpenStudy (anonymous):

@vocaloid @Tom_Boy_Rebel

OpenStudy (anonymous):

both of those people are offline XD

OpenStudy (anonymous):

Poop

OpenStudy (anonymous):

@MissNoName

OpenStudy (anonymous):

...what do I do?! I am starting to panic...I really need help to pass this

OpenStudy (anonymous):

Try C IDK!!?!

OpenStudy (anonymous):

:( you sure its c?

OpenStudy (anonymous):

Trust me try c

OpenStudy (anonymous):

@kai47 are you going to help me? lol I just got that tinge of hope XD

OpenStudy (anonymous):

ok...I put c

OpenStudy (anonymous):

-1+5=4 -1-5=-6

OpenStudy (anonymous):

do have another question now

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

man im too late

OpenStudy (anonymous):

I still have more

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

Next

OpenStudy (anonymous):

oh i can help

OpenStudy (anonymous):

YAY!

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? 2x-y=-5 x+3y=22 A. (1, 7) B. (4, 9) C. (7, 19) D. (8, 21)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A

OpenStudy (anonymous):

ok, next one?

OpenStudy (anonymous):

yes next one

OpenStudy (anonymous):

k, just a sec

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x+3y=12 -4x+3y=-3 A. (0, 4) B. (6, 2) C. (3, 3) D. (0, −1)

OpenStudy (anonymous):

...any ideas...XD

OpenStudy (anonymous):

Ooh i know

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

yes what...XD

OpenStudy (anonymous):

C

OpenStudy (anonymous):

just a sec im doing the equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think its C

OpenStudy (anonymous):

I agree C

OpenStudy (anonymous):

c looks right

OpenStudy (anonymous):

O_O the day has come...where everyone agrees on the same answer...

OpenStudy (anonymous):

XD next question? I have two more

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You can just find where they intersect on a graphing calculator.

OpenStudy (anonymous):

Tickets to the museum cost $18 for adults and $12 for children. A group of 20 people went to the museum, and the tickets cost $324. The system of equations models this situation, where x is the number of adults and y is the number of children. How many adults and how many children were in the group? x+y=20 18x+12y=324 A. 8 adults and 12 children B. 10 adults and 10 children C. 12 adults and 8 children D. 14 adults and 6 children

OpenStudy (anonymous):

i once saw a question where 4 people demanded it was an answer, one thought it was C, another thoguht D, one thought A, and one thought B, it was chaos

OpenStudy (anonymous):

O_O complete chaos...

OpenStudy (anonymous):

ickets to the museum cost $18 for adults and $12 for children. A group of 20 people went to the museum, and the tickets cost $324. The system of equations models this situation, where x is the number of adults and y is the number of children. How many adults and how many children were in the group? x+y=20 18x+12y=324 A. 8 adults and 12 children B. 10 adults and 10 children C. 12 adults and 8 children D. 14 adults and 6 children

OpenStudy (anonymous):

there...I re-posted it better XD

OpenStudy (anonymous):

Hmm, i think D

OpenStudy (anonymous):

Hold up let me think

OpenStudy (anonymous):

Yeah most likely D

OpenStudy (anonymous):

Possibly D

OpenStudy (anonymous):

D

OpenStudy (anonymous):

k...I went with d

OpenStudy (anonymous):

last one!

OpenStudy (anonymous):

Okay, lets see!

OpenStudy (anonymous):

A container holds a combination of dimes and nickels. There are 200 coins in all, and their total value is $16.30. The system of equations models this situation, where x is the number of dimes and y is the number of nickels. x+y=200 0.1x+0.05y=16.3 How many dimes and nickels are in the container? A. 126 dimes and 74 nickels B. 132 dimes and 68 nickels C. 84 dimes and 116 nickels D. 112 dimes and 88 nickels

OpenStudy (anonymous):

@vzfreakz

OpenStudy (anonymous):

okay, gimme a sec

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A? anyone else agree or disagree

OpenStudy (anonymous):

Hmm, its definitly A

OpenStudy (anonymous):

Oh!, lele yeah, its A

OpenStudy (anonymous):

Yea A

OpenStudy (anonymous):

want to know what I got...?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

oh ya, someone give kai a medal too XD

OpenStudy (anonymous):

YEA!!

OpenStudy (anonymous):

oh nvm XD

OpenStudy (anonymous):

ok...I got a...

OpenStudy (anonymous):

What'd you get?

OpenStudy (anonymous):

O_O omg...a 100%

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

100%?

OpenStudy (anonymous):

CONGRATS

OpenStudy (anonymous):

everyone...celebrate! XD

OpenStudy (anonymous):

100%!!! YES! OMG! YES!

OpenStudy (anonymous):

YAY!

OpenStudy (anonymous):

CONGRATS YOU PASSED!!

OpenStudy (anonymous):

:D I fanned all of you XD man...you guys rock XD

OpenStudy (anonymous):

Have a nice day 😊😄

OpenStudy (anonymous):

you guys too XD

OpenStudy (anonymous):

why are you guys still viewing this? XD you can fly away free now! XD

OpenStudy (anonymous):

oh sorry XD